Unproved property costs relative to the Williams lease were $40,000 at January 1, 2019. During 2019, $400,000 of drilling costs were incurred on the Williams lease. An 8%, $ 500,000 note is outstanding during the entire year and was obtained to finance the drilling program .

REQUIRED : Compute the interest capitalization amount and record the interest.
Answer #1

Interest to be capitalized = weighted average accumulated expense × interest rate on specific borrowing

First we will calculate weighted average accumulated expense:

= (40000+400000)/2 = $220000

Hence,

Interest to be capitalized = $220000 × 8% = $17600

Interest expense to be recorded = actual interest expense - expense capitalized

= ($500000×8%) - $17600.

$22400

.

i have given complete and detailed solution to your problem, in case of any query you can comment and

PLEASE DO GIVE A THUMBS UP